OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? Answer 708 412 1300 1120

OpenStudy (anonymous):

explanation?

OpenStudy (anonymous):

i would love to know where this tortured word problem comes from, because without exaggeration i have seen this here more that 20 times

OpenStudy (anonymous):

put \(x\) as the number of one step equations, and \(y\) as the number of two step equations (whatever that means) then you know \[x+y=1120\] and also \[3x-2y=1300\]

OpenStudy (tennistar):

s = one step b= two step s+b=1120 3s-2b=1300

OpenStudy (anonymous):

first equation tells us that \(y=1120-x\) so we can substitute that into the second equation for \(y\) and solve \[3x-2(1120-x)=1300\]

OpenStudy (tennistar):

Then substitute the options for these and you should find an answer

OpenStudy (anonymous):

\[3x-2240+2x=1300\] \[5x-2240=1300\] \[5x=1300+2240=3540\] \[x=3540\div 5\]

OpenStudy (tennistar):

i mean 412

OpenStudy (tennistar):

412 is 2 step 708 is one step

OpenStudy (anonymous):

i get \[x=708\] and so \[y=1120-708=412\]

OpenStudy (anonymous):

what tennistar said

OpenStudy (anonymous):

still would like to know where this comes from

OpenStudy (anonymous):

This problem comes from a Florida/Broward Virtual School Algebra 1 Course

OpenStudy (anonymous):

708 and 412